# MIRACLE II: Unveiling the multi-phase gas interplay in the circumnuclear region of NGC 1365 via multi-cloud modeling

### v3

@@ A.2 MUSE Wide Field Mode
The MUSE integrated spectra of NGC 1365 extracted from the 0.7″ radius regions shown in the lower-left panel of Fig. 1 are reported in Fig. 9. The green (blue) spectrum is extracted from a region where the outflow (stellar disk) emission is predominant. We detect approximately 30 emission lines (see Table 3) tracing different gas phases, including recombination lines, atomic lines, and lines from ionized gas.
@@ A.3 ALMA
The data include observations with two different configurations with angular resolutions in FWHM of ″ and ″, respectively. In this work, we use the observations with resolution ″ corresponding to 30 pc, which better match the MIRI MRS spatial resolution and provide a maximum recoverable scale of 2.9″. We retrieved the calibrated measurement sets using the dedicated service provided by the European ALMA Regional Center (Hatziminaoglou et al. 2015). Imaging was performed using the Common Astronomy Software Applications (CASA) v6.1.1 (CASA Team et al. 2022). For the CO(3–2) spectral window ( GHz), we subtracted a constant continuum level estimated from the emission line-free channels in the uv plane. The data were cleaned using the tclean task in CASA, with briggs weighting scheme with a robust parameter of 0.5, which resulted in a beam with a FWHM of 0.35″ 0.31″ (beam PA = -71∘), corresponding to 37 pc in physical scale. The final datacube has a spectral channel width of 8 km s-1, a pixel size of 0.05″, a FoV of 25″ in diameter and rms of 1.5 mJy beam-1 per channel.
@@ Appendix B MIRI datacubes corrections
For the astrometric registration of the 12 datacubes, we applied separate corrections for each band (SHORT, MEDIUM, and LONG), as each observing setup corresponds to a different grating configuration of the MRS. As discussed in Appendix A.1, the observations are centered on the SE outflow cone rather than the AGN. This positioning causes the source to fall near the edge, or even slightly outside, of the field of view in the Ch1 datacubes (see Fig. 1). To ensure accurate alignment across the dataset, we computed the astrometric corrections using Ch3, where the source is well within the FoV and the spatial resolution remains high, providing a reliable reference frame for registration. We collapsed the datacubes along the spectral dimension and fitted the PSFs with a 2D Gaussian model. Then, we aligned the datacubes to match the PSF peaks with the AGN coordinates provided ALMA data (see Appendix A.3). The applied corrections were RA and Dec shifts of (+0.0086″, -0.2655″), (+0.0256″, -0.2713″), and (-0.0300″, -0.299″) for the SHORT, MEDIUM, and LONG channels, respectively.
@@
As shown in Law et al. (2023), the average FWHM of the MIRI/MRS PSF varies across the four channels spanning the range FWHM = 0.4″, 0.5″, 0.6″, 0.9″, in Ch1, Ch2, Ch3, Ch4, respectively. This variation must be taken into account when computing ratios between emission lines at different wavelengths. To ensure consistent spatial resolution, we firstly downloaded the MIRI MRS PSF model cube using the WebbPSF tool v.1.4.0 (Perrin et al. 2014), which provide a model cube spatially oversampled of a factor of 4 with respect to the observations (i.e. 0.05 ″/pixel for Ch3). Then we deconvolved each flux map with the PSF corresponding to the wavelength of the emission line, and convolved it with the PSF at the reddest wavelength among the lines involved in the ratios, using the PSF model cubes from WebbPSF. The deconvolution procedure is composed by a wavelength-dependent alignment, rotation and flux normalization between the PSFs and our flux maps, following a similar methodology of Marconcini et al. (2025a).
@@ C.1 JWST/MIRI MRS emission line fitting
and the integrated flux within the FoV of the respective channel, expressed in units of erg s-1 cm-2.
[…]
We then performed the emission line fitting spaxel-by-spaxel, fitting each line in Table LABEL:tab:_table_list_emission_lines independently. Prior to this, we modeled and fitted the local continuum selecting wavelength windows on either side of the line, adjusting the range based on the FWHM of each emission line and carefully avoiding other spectral absorption or emission features. Given the short spectral window of 0.1 m we considered, within which the continuum is featureless, we fitted the continuum with a straight line. Each emission line is modeled with both a single and a double Gaussian component, and then we applied a Kolmogorov-Smirnov (KS) test on the residuals to choose the optimal number of Gaussian components in each spaxel, adopting a threshold p-value of 0.7 (see Marasco et al. 2020; Tozzi et al. 2021; Ceci et al. 2025 for details).
In the mid-IR, dust attenuation is much lower than at optical wavelengths, but it does vary with wavelength and affect emission lines differently, potentially impacting specific line ratios. Unfortunately, in our data the S/N of the Pf and Hu (i.e. H I (6-5) at 7.46 m and H I (7-6) at 12.37 m, respectively) is too low to derive a reliable extinction map directly from the mid-IR spectra. To test the potential impact of dust attenuation, we applied an extinction correction to the mid-IR emission lines using the G23 (Gordon et al. 2009, 2021, 2023; Fitzpatrick et al. 2019; Decleir et al. 2022) attenuation curve, from the dustextinction package777https://github.com/karllark/dustextinction (Gordon 2024), and adopting the maximum value (3) derived from our best-fit photoionization model (see Section 3.5.3). However, we find that the resulting variations in the mid-IR line ratios are within the errors, including those involving emission lines at distant wavelengths (e.g., [Ne V]/[Ne V], see Section 3.4).
@@ C.2 MUSE WFM emission line fitting
MUSE data were analyzed as already described in previous works of the MAGNUM program (Venturi et al. 2018; Mingozzi et al. 2019). Briefly, after a preliminary Voronoi tessellation (Cappellari & Copin 2003) to achieve an average S/N per wavelength channel of at least 70 on the continuum, we applied the penalized PiXel-Fitting (pPXF, Cappellari & Emsellem 2004) code to measure the optical emission lines listed in Table 3. Using the penalized PiXel-Fitting (pPXF, Cappellari & Emsellem 2004) code, we fitted simultaneously the continuum and gas emission in the wavelength range 4800-9200 Å, convolving a linear combination of stellar templates (E-MILES; for details see Vazdekis et al. 2016) with a Gaussian velocity distribution. Each emission line was fitted twice, using one and two Gaussian components, and the KS test was applied spaxel-by-spaxel to determine the minimum number of components required to adequately reproduce the line profiles (see Appendix C.1). Given that NGC 1365 hosts a Seyfert 1.8 nucleus, we included an additional spectral component when fitting the bins within a radius of 10 spaxels from the nuclear position to account for the spatially unresolved broad-line region (BLR). The BLR model consists of two independent components: a broad (¿1000 km s-1) Gaussian component for the broad Balmer lines (i.e. H and H) and a template of broad [Fe II] emission features by Boroson & Green (1992). The nuclear model spectrum was obtained by fitting the integrated spectrum of the central spaxels and, in the subsequent spaxel-by-spaxel fitting, it was kept fixed in shape and velocity while being allowed to scale in flux. We then subtracted the stellar and nuclear emission from the original unbinned data cube by rescaling the fitted continuum - assumed constant within each Voronoi bin - to match the observed continuum flux in each spaxel, before performing the subtraction.
@@
Finally, we corrected the optical emission-line fluxes for dust attenuation using the Balmer decrement (i.e., H/H), assuming a case-B hydrogen recombination value of 2.86 for this ratio (for an electron temperature and density of Te = K and ne = cm-3, respectively; Osterbrock & Ferland 2006) and adopting the G23 attenuation law. The results, presented in Appendix D, are consistent with previous estimates by Venturi et al. (2018) and Mingozzi et al. (2019). As shown in Fig. 12, the dust attenuation peaks at AV = 3.7 in the circumnuclear ring, where young stellar clusters are embedded (Liu et al. 2023; Schinnerer et al. 2023). High attenuation values are also found at the nuclear position, while we do not find any correlation of AV with the outflow morphology, indicative of its low dust content.
@@ Appendix D Extinction and electron density from MUSE
In this appendix, we present the spatially resolved maps of visual extinction and electron density in the circumnuclear region of NGC 1365, derived from the MUSE data and zoomed-in FoV matching the MIRI one. In the top panel of Fig. 12, we show the visual extinction computed from the Balmer decrement (H/H), adopting the G23 extinction law. High visual extinction values of A 3.7 mag are found at the nuclear position and in the southeastern part of the FoV, corresponding to the circumnuclear ring, as also observed by Venturi et al. (2018) and Mingozzi et al. (2019).
Moreover, the bottom panel of Fig. 12 shows the electron density of the ionized gas, estimated from the total intensity of all the Gaussian components of the [S II] doublet. In particular, we use the [S II]/[S II] diagnostic line ratio (Osterbrock & Ferland 2006), computed over the same FoV of the Ch4. We assume a typical value for the temperature of ionized gas of Te = 104 K and obtain the density map reported in Fig. 12. We observe that the higher values of ne are following the outflow morphology, as traced by [Ne V] contours flux. In the main text, we use as average density the median value and the standard deviation of ne = (230 120) cm-3 measured across the reported FoV.
@@ Appendix E Density comparison with IRS observations
In this appendix, we compare our spatially resolved JWST/MIRI density measurements with those obtained from integrated spectra by Dudik et al. (2007). Dudik et al. (2007) used the Infrared Spectrograph (IRS) on board Spitzer to compute the [Ne V] ratio in a sample of local AGN. They found that Type 2 AGN show higher line ratios with respect to Type 1, with all Type 2 sources lying above the theoretical LDL threshold, implying log(ne/cm-3)2. In particular, for NGC 1365 they computed a [Ne V]/[Ne V] ratio of 2.40.6, which is larger than our estimation of 1.00.2 obtained as the ratio between the two integrated fluxes in the MIRI FoV. They argued that the [Ne V] emitting region likely originates inside the nuclear torus and that differential infrared extinction due to dust in the obscuring torus may be responsible for the ratio above the LDL. Exploiting the spatially resolved [Ne V] emission in our data and the density values in the nuclear region, we can rule out this interpretation. However, given the presence of some LDL regions in Fig. 6, it is possible that their finding is the result of spatial integration over a wide slit, thus including such regions in the measurement. A more plausible reason is an observational effect. Indeed, in Dudik et al. (2007) the [Ne V] fluxes were obtained using the short-wavelength, high-resolution (SH; 4.7″ 11.3″, = 9.9-19.6 m) and long-wavelength, high-resolution (LH; 11.1″ 22.3″, = 18.7-37.2 m) modules of IRS. As it is clear from the moment maps of the high-IP species in the online material, the extension of [Ne V] emission is detected up to 8″ from the nucleus. This means that the SH module surely lost part of the [Ne V] flux, leading to an artificially enhancement of the [Ne V]/[Ne V] line ratio.
@@ Appendix G Ionized gas mass estimate
We assume , 1.2 (i.e. a 10% number density of He atoms with respect to H atoms), and a solar abundance of [Ne/H] 7.93 (Asplund et al. 2009), to obtain . Assuming a typical temperature (T), from the [Ne V]/[Ne V] line ratio we find an electron density of ne = (750440) cm-3 in the 1.5” radius aperture and estimate a line emissivity of erg s-1 cm3 making use of PyNeb (Luridiana et al. 2015). Finally, Eq. 6 can be rewritten as
@@
Similarly, we compute the [O III] outflow mass. Using a solar abundance of [O/H] 8.69 (Asplund et al. 2009) and ne = (500 270) cm-3 (from the [S II] ratio in the same aperture) - that leads to erg s-1 cm3 - we have
We corrected both the [Ne V] and the [O III] luminosity for dust attenuation using the Balmer decrement from the Hα/Hβ ratio, employing the G23 attenuation law and an intrinsic ratio Hα/Hβ = 2.86 (for Te = 104 K, Osterbrock & Ferland 2006). Assuming the two different densities for the optical and mid-IR species previously computed, we obtain an outflow mass of = (2.6 1.0) 103 and = (20 7) 104 .

### v4

@@ A.2 MUSE Wide Field Mode
The MUSE integrated spectra of NGC 1365 extracted from the 0.7″ radius regions shown in the lower left panel of Fig. 1 are reported in Fig. 9. The green (blue) spectrum is extracted from a region where the outflow (stellar disk) emission is predominant. We detect approximately 30 emission lines (see Table 3) tracing different gas phases, including recombination lines, atomic lines, and lines from ionized gas.
@@ A.3 ALMA
The data include observations with two different configurations with angular resolutions in FWHM of ″ and ″, respectively. In this work, we use the observations with resolution ″ corresponding to 30 pc, which better match the MIRI MRS spatial resolution and provide a maximum recoverable scale of 2.9″. We retrieved the calibrated measurement sets using the dedicated service provided by the European ALMA Regional Center (Hatziminaoglou et al. 2015). Imaging was performed using the Common Astronomy Software Applications (CASA) v6.1.1 (CASA Team et al. 2022). For the CO(3–2) spectral window ( GHz), we subtracted a constant continuum level estimated from the emission line-free channels in the uv plane. The data were cleaned using the tclean task in CASA, with briggs weighting scheme with a robust parameter of 0.5, which resulted in a beam with a FWHM of 0.35″ 0.31″ (beam PA = -71∘), corresponding to 37 pc in physical scale. The final datacube has a spectral channel width of 8 km s-1, a pixel size of 0.05″, a FoV of 25″ in diameter, and a root mean square of 1.5 mJy beam-1 per channel.
@@ Appendix B MIRI datacubes corrections
For the astrometric registration of the 12 datacubes, we applied separate corrections for each band (SHORT, MEDIUM, and LONG), as each observing setup corresponds to a different grating configuration of the MRS. As is discussed in Appendix A.1, the observations are centered on the SE outflow cone rather than the AGN. This positioning causes the source to fall near the edge, or even slightly outside, of the FoV in the Ch1 datacubes (see Fig. 1). To ensure accurate alignment across the dataset, we computed the astrometric corrections using Ch3, where the source is well within the FoV and the spatial resolution remains high, providing a reliable reference frame for registration. We collapsed the datacubes along the spectral dimension and fit the PSFs with a 2D Gaussian model. Then, we aligned the datacubes to match the PSF peaks with the AGN coordinates provided ALMA data (see Appendix A.3). The applied corrections were RA and Dec shifts of (+0.0086″, -0.2655″), (+0.0256″, -0.2713″), and (-0.0300″, -0.299″) for the SHORT, MEDIUM, and LONG channels, respectively.
@@
As shown in Law et al. (2023), the average FWHM of the MIRI/MRS PSF varies across the four channels spanning the range FWHM = 0.4″, 0.5″, 0.6″, and 0.9″, in Ch1, Ch2, Ch3, and Ch4, respectively. This variation must be taken into account when computing ratios between emission lines at different wavelengths. To ensure consistent spatial resolution, we firstly downloaded the MIRI MRS PSF model cube using the WebbPSF tool v.1.4.0 (Perrin et al. 2014), which provide a model cube spatially oversampled of a factor of 4 with respect to the observations (i.e., 0.05 ″/pixel for Ch3). Then we deconvolved each flux map with the PSF corresponding to the wavelength of the emission line, and convolved it with the PSF at the reddest wavelength among the lines involved in the ratios, using the PSF model cubes from WebbPSF. The deconvolution procedure is composed by a wavelength-dependent alignment, rotation and flux normalization between the PSFs and our flux maps, following a similar methodology of Marconcini et al. (2025a).
@@ C.1 JWST/MIRI MRS emission line fitting
(1) ISO Spectrometer Data Centre.
We then performed the emission line fitting spaxel-by-spaxel, fitting each line in Table LABEL:tab:_table_list_emission_lines independently. Prior to this, we modeled and fit the local continuum selecting wavelength windows on either side of the line, adjusting the range based on the FWHM of each emission line and carefully avoiding other spectral absorption or emission features. Given the short spectral window of 0.1 m we considered, within which the continuum is featureless, we fit the continuum with a straight line. Each emission line is modeled with both a single and a double Gaussian component, and then we applied a Kolmogorov-Smirnov (KS) test on the residuals to choose the optimal number of Gaussian components in each spaxel, adopting a threshold p-value of 0.7 (see Marasco et al. 2020; Tozzi et al. 2021; Ceci et al. 2025 for details).
In the mid-IR, dust attenuation is much lower than at optical wavelengths, but it does vary with wavelength and affect emission lines differently, potentially impacting specific line ratios. Unfortunately, in our data the S/N of the Pf and Hu (i.e., H I (6-5) at 7.46 m and H I (7-6) at 12.37 m, respectively) is too low to derive a reliable extinction map directly from the mid-IR spectra. To test the potential impact of dust attenuation, we applied an extinction correction to the mid-IR emission lines using the G23 (Gordon et al. 2009, 2021, 2023; Fitzpatrick et al. 2019; Decleir et al. 2022) attenuation curve, from the dustextinction package777https://github.com/karllark/dustextinction (Gordon 2024), and adopting the maximum value (3) derived from our best-fit photoionization model (see Section 3.5.3). However, we find that the resulting variations in the mid-IR line ratios are within the errors, including those involving emission lines at distant wavelengths (e.g., [Ne V]/[Ne V], see Section 3.4).
@@ C.2 MUSE WFM emission line fitting
MUSE data were analyzed as already described in previous works of the MAGNUM program (Venturi et al. 2018; Mingozzi et al. 2019). Briefly, after a preliminary Voronoi tessellation (Cappellari and Copin 2003) to achieve an average S/N per wavelength channel of at least 70 on the continuum, we applied the penalized PiXel-Fitting (pPXF, Cappellari and Emsellem 2004) code to measure the optical emission lines listed in Table 3. Using the penalized PiXel-Fitting (pPXF, Cappellari and Emsellem 2004) code, we fit simultaneously the continuum and gas emission in the wavelength range 4800-9200 Å, convolving a linear combination of stellar templates (E-MILES; for details see Vazdekis et al. 2016) with a Gaussian velocity distribution. Each emission line was fit twice, using one and two Gaussian components, and the KS test was applied spaxel-by-spaxel to determine the minimum number of components required to adequately reproduce the line profiles (see Appendix C.1). Given that NGC 1365 hosts a Seyfert 1.8 nucleus, we included an additional spectral component when fitting the bins within a radius of 10 spaxels from the nuclear position to account for the spatially unresolved broad-line region (BLR). The BLR model consists of two independent components: a broad (¿1000 km s-1) Gaussian component for the broad Balmer lines (i.e., H and H) and a template of broad [Fe II] emission features by Boroson and Green (1992). The nuclear model spectrum was obtained by fitting the integrated spectrum of the central spaxels and, in the subsequent spaxel-by-spaxel fitting, it was kept fixed in shape and velocity while being allowed to scale in flux. We then subtracted the stellar and nuclear emission from the original unbinned data cube by rescaling the fit continuum - assumed constant within each Voronoi bin - to match the observed continuum flux in each spaxel, before performing the subtraction.
@@
Finally, we corrected the optical emission-line fluxes for dust attenuation using the Balmer decrement (i.e., H/H), assuming a case-B hydrogen recombination value of 2.86 for this ratio (for an electron temperature and density of Te = K and ne = cm-3, respectively; Osterbrock and Ferland 2006) and adopting the G23 attenuation law. The results, presented in Appendix D, are consistent with previous estimates by Venturi et al. (2018) and Mingozzi et al. (2019). As shown in Fig. 12, the dust attenuation peaks at AV = 3.7 in the circumnuclear ring, where young stellar clusters are embedded (Liu et al. 2023; Schinnerer et al. 2023). High attenuation values are also found at the nuclear position, while we do not find any correlation of AV with the outflow morphology, indicative of its low dust content.
@@ Appendix D Extinction and electron density from MUSE
In this appendix, we present the spatially resolved maps of visual extinction and electron density in the circumnuclear region of NGC 1365, derived from the MUSE data and zoomed-in FoV matching the MIRI one. In the top panel of Fig. 12, we show the visual extinction computed from the Balmer decrement (H/H), adopting the G23 extinction law. High visual extinction values of A 3.7 mag are found at the nuclear position and in the SE part of the FoV, corresponding to the circumnuclear ring, as also observed by Venturi et al. (2018) and Mingozzi et al. (2019).
Moreover, the bottom panel of Fig. 12 shows the electron density of the ionized gas, estimated from the total intensity of all the Gaussian components of the [S II] doublet. In particular, we use the [S II]/[S II] diagnostic line ratio (Osterbrock and Ferland 2006), computed over the same FoV of the Ch4. We assume a typical value for the temperature of ionized gas of Te = 104 K and obtain the density map reported in Fig. 12. We observe that the higher values of ne are following the outflow morphology, as traced by [Ne V] contours flux. In the main text, we use as average density the median value and the standard deviation of ne = (230 120) cm-3 measured across the reported FoV.
@@ Appendix E Density comparison with IRS observations
In this appendix, we compare our spatially resolved JWST/MIRI density measurements with those obtained from integrated spectra by Dudik et al. (2007). Dudik et al. (2007) used the Infrared Spectrograph (IRS) on board Spitzer to compute the [Ne V] ratio in a sample of local AGNs. They found that Type 2 AGNs show higher line ratios with respect to Type 1, with all Type 2 sources lying above the theoretical LDL threshold, implying log(ne/cm-3)2. In particular, for NGC 1365 they computed a [Ne V]/[Ne V] ratio of 2.40.6, which is larger than our estimation of 1.00.2 obtained as the ratio between the two integrated fluxes in the MIRI FoV. They argued that the [Ne V] emitting region likely originates inside the nuclear torus and that differential IR extinction due to dust in the obscuring torus may be responsible for the ratio above the LDL. Exploiting the spatially resolved [Ne V] emission in our data and the density values in the nuclear region, we can rule out this interpretation. However, given the presence of some LDL regions in Fig. 6, it is possible that their finding is the result of spatial integration over a wide slit, and thus including such regions in the measurement. A more plausible reason is an observational effect. Indeed, in Dudik et al. (2007) the [Ne V] fluxes were obtained using the short-wavelength, high-resolution (SH; 4.7″ 11.3″, = 9.9-19.6 m) and long-wavelength, high-resolution (LH; 11.1″ 22.3″, = 18.7-37.2 m) modules of IRS. As it is clear from the moment maps of the high-IP species in the online material, the extension of [Ne V] emission is detected up to 8″ from the nucleus. This means that the SH module surely lost part of the [Ne V] flux, leading to an artificially enhancement of the [Ne V]/[Ne V] line ratio.
@@ Appendix G Ionized gas mass estimate
We assume , 1.2 (i.e., a 10% number density of He atoms with respect to H atoms), and a solar abundance of [Ne/H] 7.93 (Asplund et al. 2009), to obtain . Assuming a typical temperature (T), from the [Ne V]/[Ne V] line ratio we find an electron density of ne = (750440) cm-3 in the 1.5” radius aperture and estimate a line emissivity of erg s-1 cm3 making use of PyNeb (Luridiana et al. 2015). Finally, Eq. 6 can be rewritten as
@@
Similarly, we computed the [O III] outflow mass. Using a solar abundance of [O/H] 8.69 (Asplund et al. 2009) and ne = (500 270) cm-3 (from the [S II] ratio in the same aperture) - that leads to erg s-1 cm3 - we have
We corrected both the [Ne V] and the [O III] luminosity for dust attenuation using the Balmer decrement from the Hα/Hβ ratio, employing the G23 attenuation law and an intrinsic ratio Hα/Hβ = 2.86 (for Te = 104 K, Osterbrock and Ferland 2006). Assuming the two different densities for the optical and mid-IR species previously computed, we obtain an outflow mass of = (2.6 1.0) 103 and = (20 7) 104 .